# This is a preprint submitted to: Archives of Computational Methods in Engineering

### v1

@@ 4.2 Geometric filtering
Several case studies where geometric filtering has been applied have been found in literature, specially for aerodynamics problems. An engineering-knowledge-based filtering model using support vector regression to enhance the optimization of a two-dimensional intake duct by incorporating several geometric and flow parameters was developed in [22]. An RBF network to eliminate abnormal computer-aided design (CAD) models in nacelle design is employed in [23]. An inverse distance weighting to define geometric filtering constraints in modal parameterization of airfoils, which helped train more accurate data-based airfoil analysis models is used in [24]. Using deep learning, a validity model is created in [25] to detect abnormal geometric shapes in airfoils and wing sections, which was trained on a large dataset of labeled airfoils and used in conjunction with generative adversarial networks (GANs) to improve the realism of synthetic airfoils. The effectiveness of geometric filtering in maintaining innovative designs that enhance aerodynamic performance is also noted, with findings supporting its utility in optimizing designs within transonic regimes [26].
@@ 5 Indirect Dimensionality Reduction Methods
Indirect dimensionality reduction methods do not alter the data space itself but focus on identifying the most influential variables and constraints.
@@
Once the influential factors are identified through this combined screening approach, incorporating both the straightforward yet powerful Morris method and other techniques, further optimization efforts can be concentrated on adjusting these key factors to achieve the desired performance goals or characteristics of the optimized shape. This iterative process helps to streamline the optimization process and improve the overall efficiency of shape optimization studies. However, this approach does not always provide the best solution, as it does not fully evaluate the importance that the fixed variables could have during the optimization process, especially when combined with other variables. This limitation can be assessed by variance-based methods like Sobol sensitivity analysis [29]. The preliminary design of a bulk carrier to enhance its robustness and reliability by focusing only on the most influential random variables, thereby reducing the complexity of the optimization problem, was performed in [30], where polynomial chaos expansion was employed in combination with sensitive analysis to calculate Sobol’s indices, by deriving them directly from the coefficients of the polynomial representing the output variable. The approximate local solution of Sobol’s indices was applied to the reliability-based robust design optimization of composite structures in [31]. Nevertheless, the Sobol method operates within a probabilistic framework and consequently requires a statistically significant number of design-space samples. Furthermore, the number of indices increases with the power of the design-space dimensionality, making Sobol indices generally computationally expensive.
@@ 6 Direct Dimensionality Reduction Methods: Representation Learning
$\min_{{\mathbf{x}}}\quad$ $f({\mathbf{x}},{\mathbf{y}})$ (3)
subject to $g_{i}({\mathbf{x}},{\mathbf{y}})\leq 0,\quad i=1,\ldots,I$
and to $h_{j}({\mathbf{x}},{\mathbf{y}})=0,\quad j=1,\ldots,J$
@@
Figure 12: Simulation-based design optimization XDSM diagram for shape optimization extended with dimensionality reduction based on PCA
@@ 6.1 Linear dimensionality reduction methods
6.1.1 Proper orthogonal decomposition/Karhunen-Loève expansion leading to principal component analysis: from continuum to discrete geometry-based formulation
@@
Discretizing the domain $\mathcal{G}$ into $L$ elements, each with a measure denoted as $\Delta\mathcal{G}_{i}$ (where $i=1,\dots,L$), and sampling the input space $\mathcal{U}$ using a statistically convergent number of Monte Carlo (MC) realizations, denoted as $S$, results in a set $\{{\mathbf{u}}_{k}\}_{k=1}^{S}$ drawn from the distribution $p({\mathbf{u}})$. These realizations are organized into a data matrix $\mathbf{D}$ of dimensionality $\left[3L\times S\right]$, given by:
@@
Figure 13: Example of (left) MC samples of the design space for the optimization of a hydrofoil [34] and (right) dimensionality reduction results in terms of cumulative geometric variance retained by the PCA components
@@
The XDSM diagram shown in Fig. 2 can be consequently transformed as shown in Fig. 12
POD/PCA have been crucial in enhancing the design of airfoils and wings, streamlining external aerodynamic optimizations for both subsonic [35, 36, 24] and transonic [37, 38, 39, 40, 41, 42, 24, 43, 44, 45, 46] regimes of airfoils and wings, as well as for internal aerodynamics of compressors [47, 48], turbines [49], and nozzles [50]. These methods have facilitated efficient parameterization and reduced drag, optimizing aerodynamic performance for aircraft, including underwater autonomous vehicles [51]. In marine applications, KLE/PCA has improved hydrodynamic performance by optimizing the designs of mono- [4, 52, 53, 54] and multi-hulls [33, 55, 56], hull vanes [57], and propellers [58]. These applications have demonstrated effective reductions in design variables and computational costs while improving resistance and efficiency. Naval hull optimizations [59] have particularly benefited from generative models and advanced PCA techniques [60] that manage design deformations and enhance performance. Figure 13 shows an example of MC samples (used for the construction of the data matrix $\mathbf{D}$) of $M=10$ parametric design space for the optimization of a hydrofoil and the associated dimensionality reduction result [34].
@@ 6.1.2 Physics-informed principal component analysis
Design-space dimensionality reduction using PCA traditionally relies solely on geometric information. However, by augmenting the data matrix $\mathbf{D}$ with physical data derived from numerical simulations, the method’s applicability and effectiveness can be substantially enhance. This augmentation can include both distributed values, such as velocity or pressure fields, and lumped parameters like, e.g., resistance or efficiency. By integrating these additional data types, the PCA not only captures geometric variability but also reflects critical physical phenomena that could influence the design’s performance.
Incorporating physics-informed parameters allows for a more holistic understanding of the design space by bridging the gap between purely geometric insights and their functional outcomes. For example, the use of PCA in conjunction with physics-informed data to analyze ship hydrodynamics, assessing how variations in shape impact physical parameters like wave resistance, has been in demonstrated in [61]. This approach aligns with the work in [62], where PCA extended with simulation data effectively mapped the design space of waterjet propulsion systems, considering not only the shape but also the flow characteristics essential for optimizing performance. Furthermore, [63, 64] explored a combined shape- and physics-based PCA formulation, which significantly improved the efficiency of hydrodynamic shape optimizations by including low-fidelity solver outputs in the preliminary design phases. This method efficiently identifies key design modifications that contribute to enhanced deterministic [64] and stochastic [63] performance, thus streamlining the optimization process under multi-fidelity settings. Additionally, [65, 66, 67] introduced geometric moments as cost-effective, physics-related descriptors that substitute direct simulation data, further enriching the PCA model without the computational cost of numerical simulations. This adaptation enables the efficient exploration of design spaces by focusing on features that are both geometrically significant and closely related to performance metrics.
These enhancements to PCA not only increase the accuracy of the dimensionality reduction but also enable designers to make more informed decisions early in the design process, potentially reducing the number of high-fidelity simulations required and thereby lowering computational costs.
@@ 6.1.3 Parametric model embedding
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Having given a null weight to $\mathbf{U}$ does not remove the contribution of the design variables from the inner product, but just cancels as many columns as $M$ from the matrix $\widetilde{\mathbf{A}}\widetilde{\mathbf{G}}\widetilde{\mathbf{W}}$, thus Eqs. 15 and 23 provides the same eigenvalues ($\boldsymbol{\Lambda}=\widetilde{\boldsymbol{\Lambda}}$) and geometrical components of the eigenvectors (${\mathbf{z}}_{k}$, except for a multiplicative constant). The proof of the equivalence between KLE/PCA and PME is given in [71]. In addition and as desired, the solution of Eq. 23 provides the eigenvector components ${\mathbf{v}}_{k}$ that embed the original design variables ${\mathbf{u}}$ (see Fig. 14).
@@
where the components of the eigenvector ${\mathbf{v}}_{k}$ embed (or contain) the reduced-order representation of the original design parameterization.
The method has been presented in [71] and successfully applied for both the aerodynamics optimization of NACA and RAE [68] foils and the hydrodynamic optimization of a destroyer-type vessel.
@@ 6.1.4 Active subspace method
Consider the objective function $f$ of problem 2 that models the system of interest. Assuming $f$ as differentiable with respect to the inputs ${\mathbf{u}}$, the gradient of $f$, denoted by $\nabla f({\mathbf{u}})=\left[\frac{\partial f}{\partial u_{1}}\,\,\dots\,\,\frac{\partial f}{\partial u_{M}}\right]^{\mathsf{T}}$, contains information about the sensitivity of the output with respect to changes in each input variable.
@@
This transformation enables the exploration of the most influential directions with fewer variables (see Fig. 15), significantly reducing the computational complexity of the optimization process.
@@
ASM has been effectively used to reduce high-dimensional design spaces in external and internal [73, 74] aerodynamic optimizations, facilitating the exploration and improvement of airfoil [75, 76] and wing [77, 78], aero-engine nacelle [79, 80], as well as satellite applications [81]. This has helped in achieving better performance metrics like lift-to-drag ratios by focusing on the most influential variables, leading to computationally efficient designs. In naval engineering, ASM has supported the optimization of ship hulls and other marine structures, simplifying the complex sets of design variables into manageable subspaces and thus improving hydrodynamic performance and structural integrity [82, 83].
@@ 6.2 Nonlinear dimensionality reduction methods
Figure 16: Example of LPCA analysis with first components
@@ 6.2.1 Local principal component analysis
LPCA is particularly useful in shape optimization where the design space is inherently complex and multidimensional. By localizing the PCA, LPCA can adapt more flexibly to local variations within the data, capturing essential features that global PCA might miss. Each local model can offer insights into specific regions of the design space, potentially revealing localized trends and dependencies that are relevant for detailed design analysis.
@@ 6.2.2 Kernel principal component analysis
Then PCA is computed in the feature space $\mathcal{F}$. Assuming that $\sum_{k}\Gamma(\mathbf{d}({\mathbf{u}}_{k}))=0$, the kernel principal component $[\mathbf{z}_{1}\,\dots\,{\mathbf{z}}_{J}]$ can be find solving the eigenproblem
@@
The reconstruction of the original data from the feature space $\mathcal{F}$ in KPCA is more problematic than PCA, since it needs to find, for every point $\Gamma(\mathbf{d}({\mathbf{u}}_{k}))$, the relative pre-image $\mathbf{d}({\mathbf{u}}_{k})$ in the input space $\mathcal{I}$. Nevertheless, KPCA has been successfully applied for the design-space assessment of a destroyer-type vessel in [88], where the approximate pre-images technique proposed in [90] is used, and aerodynamic optimization of NACA foil in [69].

### v2

@@ 4.2 Geometric filtering
Several case studies where geometric filtering has been applied have been found in the literature. An engineering-knowledge-based filtering model using support vector regression to enhance the optimization of a two-dimensional intake duct by incorporating several geometric and flow parameters was developed in [35]. An RBF network to eliminate abnormal computer-aided design (CAD) models in nacelle design is employed in [36]. An inverse distance weighting to define geometric filtering constraints in modal parameterization of airfoils, which helped train more accurate data-based airfoil analysis models is used in [37]. Using deep learning, a validity model is created in [38] to detect abnormal geometric shapes in airfoils and wing sections, which was trained on a large dataset of labeled airfoils and used in conjunction with generative adversarial networks (GANs) to improve the realism of synthetic airfoils. The effectiveness of geometric filtering in maintaining innovative designs that enhance aerodynamic performance is also noted, with findings supporting its utility in optimizing designs within transonic regimes [39].
@@ 5 Indirect Dimensionality Reduction Methods
Indirect dimensionality reduction methods do not alter the data space itself but focus on identifying the most influential variables and constraints (see Fig. 12).
@@
Figure 12: Block diagram for the solution of shape optimization problems through sensitivity methods
@@
Once the influential factors are identified through this combined screening approach, incorporating both the straightforward yet powerful Morris method and other techniques, further optimization efforts can be concentrated on adjusting these key factors to achieve the desired performance goals or characteristics of the optimized shape. This iterative process helps to streamline the optimization process and improve the overall efficiency of shape optimization studies. For instance, the influence of design variables on the objective function is calculated by regression models based on the radial basis functions in [42] and Kriging [43].
However, this approach does not always provide the best solution, as it does not fully evaluate the importance that the fixed variables could have during the optimization process, especially when combined with other variables. This limitation can be assessed by variance-based methods like Sobol sensitivity analysis [44, 45]. The preliminary design of a bulk carrier to enhance its robustness and reliability by focusing only on the most influential random variables, thereby reducing the complexity of the optimization problem, was performed in [46], where polynomial chaos expansion was employed in combination with sensitive analysis to calculate Sobol’s indices, by deriving them directly from the coefficients of the polynomial representing the output variable. The approximate local solution of Sobol’s indices was applied to the reliability-based robust design optimization of composite structures in [47]. An improved variance-based sensitivity analysis method was developed specifically for ship hull optimization, where the Sobol method was enhanced by introducing new integrals characterized by reduced variance and employing quasi-random numbers with lower star discrepancy [48]. This enhancement significantly improved the accuracy of sensitivity indices estimation, facilitating an effective reduction of the dimensionality of the hull-form optimization model. Nevertheless, the Sobol method operates within a probabilistic framework and consequently requires a statistically significant number of design-space samples. Furthermore, the number of indices increases with the power of the design-space dimensionality, making Sobol indices generally computationally expensive. It is worth noting that surrogate-model-based Sobol analyses can significantly mitigate this computational burden, offering more efficient estimates of sensitivity indices; however, such surrogate-based strategies were not extensively identified within the reviewed shape optimization literature.
@@ 6 Direct Dimensionality Reduction Methods: Representation Learning
$\min_{{\mathbf{x}}}\quad$ $f({\mathbf{x}})$ (3)
subject to $g_{i}({\mathbf{x}})\leq 0,\quad i=1,\ldots,I$
and to $h_{j}({\mathbf{x}})=0,\quad j=1,\ldots,J$
@@ 6.1 Linear dimensionality reduction methods
6.1.1 Principal component analysis, from KLE/POD
@@
Figure 14: Block diagram for the solution of shape optimization problems through PCA
Figure 15: Example of (left) random samples of the design space for the optimization of a hydrofoil [51] and (right) dimensionality reduction results in terms of cumulative geometric variance retained by the PCA components
Discretizing the domain $\mathcal{G}$ into $L$ elements, each with a measure denoted as $\Delta\mathcal{G}_{i}$ (where $i=1,\dots,L$), and sampling the input space $\mathcal{U}$ using a statistically convergent number of realizations, denoted as $S$, results in a set $\{{\mathbf{u}}_{k}\}_{k=1}^{S}$ drawn from the distribution $p({\mathbf{u}})$. These realizations are organized into a data matrix $\mathbf{D}$ of dimensionality $\left[3L\times S\right]$, given by:
@@
The block diagram shown in Fig. 2 can be consequently transformed as shown in Fig. 14, where it is shown how the dimensionality reduction procedure is performed upfront the optimization loop and fed it with the new parameterization based on PCA eigenvectors and eigenvalues.
POD/PCA have been crucial in enhancing the design of airfoils and wings, streamlining external aerodynamic optimizations for both subsonic [52, 53, 37], transonic [54, 55, 56, 57, 58, 59, 37, 60, 61, 62, 63, 64, 65], and hypersonic [66] regimes of airfoils and wings, as well as for internal aerodynamics of compressors [67, 68], turbines [69], and nozzles [70]. These methods have facilitated efficient parameterization and reduced drag, optimizing aerodynamic performance for aircraft and watercraft, including blended wing body [71] and underwater autonomous [72] vehicles. In marine applications, KLE/PCA has improved hydrodynamic performance by optimizing the designs of mono- [4, 73, 74, 75, 76] and multi-hulls [50, 77, 78], hull vanes [79], propellers [80], and turbines [81]. These applications have demonstrated effective reductions in design variables and computational costs while improving resistance and efficiency. Naval hull optimizations [82] have particularly benefited from generative models and advanced PCA techniques [83] that manage design deformations and enhance performance. Figure 15 shows an example of random samples (used for the construction of the data matrix $\mathbf{D}$) of $M=10$ parametric design space for the optimization of a hydrofoil and the associated dimensionality reduction result [51] in terms of geometric variance resolved as a function of the number of reduced design variables.
@@
Figure 16: Block diagram for the solution of shape optimization problems through physics-informed PCA
@@ 6.1.2 Physics-informed principal component analysis
Design-space dimensionality reduction using PCA traditionally relies solely on geometric information. However, by augmenting the data matrix $\mathbf{D}$ with physical data derived from numerical simulations, the method’s applicability and effectiveness can be substantially enhance. This augmentation can include both distributed values, such as velocity or pressure fields, and lumped parameters like, e.g., resistance or efficiency. By integrating these additional data types, the PCA not only captures geometric variability but also reflects critical physical phenomena that could influence the design’s performance. Figure 16 shows how PCA-based workflow (see Fig. 14) can be augmented by computing low-fidelity physical quantities for each design variant of the original design space upfront the optimization loop.
Incorporating physics-informed parameters allows for a more holistic understanding of the design space by bridging the gap between purely geometric insights and their functional outcomes. For example, PCA integrated with physical data has been demonstrated to effectively map the design space of waterjet propulsion systems, considering shape and flow characteristics essential for optimizing performance [84]. Similarly, aerodynamic shape optimization of turbomachinery has benefited from the integration of aerodynamic sensitivities into PCA, improving uncertainty quantification by capturing geometry variations critical to performance metrics [85, 86] and enhancing the efficiency of deviation modeling in compressors [86]. Furthermore, a metric-based POD parameterization method has demonstrated significant improvements in aerodynamic shape optimization efficiency, employing data-based filtering strategies to refine design spaces specifically towards better aerodynamic characteristics [87]. In rotorcraft applications, aerodynamic and aeroacoustic optimization has been substantially advanced by integrating POD parameterization with physical aerodynamic data, enabling accurate and efficient exploration of design spaces sensitive to performance and noise constraints [88]. In naval hydrodynamics, PCA combined with physics-informed simulation data has been successfully employed to analyze ship hydrodynamics, assessing how shape variations influence wave resistance and other critical performance parameters [89]. Moreover, recent studies explored a combined shape- and physics-based PCA formulation that significantly improves hydrodynamic shape optimization efficiency by integrating low-fidelity solver outputs into preliminary design phases, thereby efficiently identifying key design modifications contributing to enhanced deterministic [90] and stochastic performance [91] under multi-fidelity conditions.
Additionally, [92, 93, 94, 95, 96] introduced geometric moments as cost-effective, physics-related descriptors that substitute direct simulation data, further enriching the PCA model without the computational cost of numerical simulations. This adaptation enables the efficient exploration of design spaces by focusing on features that are both geometrically significant and closely related to performance metrics.
These enhancements to PCA, leveraging physics-informed methods, not only increase the accuracy of the dimensionality reduction but also enable designers to make more informed decisions early in the design process, potentially reducing the number of high-fidelity simulations required and thereby lowering computational costs.
@@ 6.1.3 Parametric model embedding
Figure 17: Block diagram for the solution of shape optimization problems through PME
@@
Having given a null weight to $\mathbf{U}$ does not remove the contribution of the design variables from the inner product, but just cancels as many columns as $M$ from the matrix $\widetilde{\mathbf{A}}\widetilde{\mathbf{G}}\widetilde{\mathbf{W}}$, thus Eqs. 15 and 23 provides the same eigenvalues ($\boldsymbol{\Lambda}=\widetilde{\boldsymbol{\Lambda}}$) and geometrical components of the eigenvectors (${\mathbf{z}}_{k}$, except for a multiplicative constant). The proof of the equivalence between KLE/PCA and PME is given in [99]. In addition and as desired, the solution of Eq. 23 provides the eigenvector components ${\mathbf{v}}_{k}$ that embed the original design variables ${\mathbf{u}}$.
@@
where the components of the eigenvector ${\mathbf{v}}_{k}$ embed (or contain) the reduced-order representation of the original design parameterization. Figure 17 shows the integration of the original design variables in the dimensionality reduction procedure and how the optimization loop is modified by retrieving the original parametrization through PME eingevectors.
The method has been presented in [99] and successfully applied for both the aerodynamics optimization of NACA and RAE [101] foils and the hydrodynamic optimization of a destroyer-type vessel [102]. Figure 18 provide an example that compares PCA and PME methods by visualizing the first three eigenvector components for both approaches applied to the shape optimization of the RAE-2822 airfoil [101]. In particular, the top row shows the shape modification associated with the PCA re-parameterization. The bottom row, instead, gives the eigenvectors ${\mathbf{v}}_{k}$ that enable the explicit mapping between reduced coordinates and original design variables. Furthermore, looking at peaks of each ${\mathbf{v}}_{k}$ components, PME allows to assess which of the original design parameters mostly influence the associated portion of geometric variance retained.
Figure 19: Block diagram for the solution of shape optimization problems through ASM
@@ 6.1.4 Active subspace method
Consider the objective function $f$ of problem 2 that models the system of interest. Assuming $f$ as differentiable with respect to the inputs ${\mathbf{u}}$, the gradient of $f$, denoted by
$\nabla f({\mathbf{u}})=\left[\frac{\partial f}{\partial u_{1}}\,\,\dots\,\,\frac{\partial f}{\partial u_{M}}\right]^{\mathsf{T}},$ (28)
contains information about the sensitivity of the output with respect to changes in each input variable.
@@
This transformation enables the exploration of the most influential directions with fewer variables, significantly reducing the computational complexity of the optimization process.
Figure 19 shows the block diagram associated with ASM dimensionality reduction procedure and subsequent optimization loop, where the original parameterization is retrieved through back-mapping as per Eq. 32. A simple example illustrating how the ASM operates on a bi-dimensional design space is shown in Fig. 20. The figure clearly shows that the gradient of the objective function predominantly follows the diagonal direction of the design space. This indicates that the function can be effectively represented in a one-dimensional active subspace, defined by a single reduced variable.
@@
ASM has been effectively used to reduce high-dimensional design spaces in external and internal [104, 105] aerodynamic optimizations, facilitating the exploration and improvement of airfoil [106, 107] and wing [108, 109, 110], aero-engine nacelle [111, 112], as well as satellite applications [113], including also fluid-structure interaction [114] and uncertainty quantification [115] problems. This has helped in achieving better performance metrics like lift-to-drag ratios by focusing on the most influential variables, leading to computationally efficient designs. In naval engineering, ASM has supported the optimization of ship hulls and other marine structures, simplifying the complex sets of design variables into manageable subspaces and thus improving hydrodynamic performance and structural integrity [116, 117]. Additionally, in the field of internal combustion engines, ASM has recently been employed to analyze and reduce the dimensionality of complex design problems related to the combustion chamber structure of Wankel rotary engines, facilitating the identification of critical structural parameters and significantly enhancing optimization efficiency [118].
@@ 6.2 Nonlinear dimensionality reduction methods
Figure 21: LPCA examples with first components
@@
Figure 23: Block diagram for the solution of shape optimization problems through LPCA
@@ 6.2.1 Local principal component analysis
LPCA is particularly useful in shape optimization where the design space is inherently complex and multidimensional. By localizing the PCA, LPCA can adapt more flexibly to local variations within the data, capturing essential features that global PCA might miss. Each local model can offer insights into specific regions of the design space, potentially revealing localized trends and dependencies that are relevant for detailed design analysis. Figure 23 depicts the block diagram using LPCA, involving partitioning the design space into distinct clusters and performing PCA locally within each cluster.
@@
Figure 24: Block diagram for the solution of shape optimization problems through KPCA
@@ 6.2.2 Kernel principal component analysis
Then PCA is computed in the feature space $\mathcal{F}$ (see Fig. 24). Assuming that $\sum_{k}\Gamma(\mathbf{d}({\mathbf{u}}_{k}))=0$, the kernel principal component $[\mathbf{z}_{1}\,\dots\,{\mathbf{z}}_{J}]$ can be find solving the eigenproblem
@@
The reconstruction of the original data from the feature space $\mathcal{F}$ in KPCA is more problematic than PCA, since it needs to find, for every point $\Gamma(\mathbf{d}({\mathbf{u}}_{k}))$, the relative pre-image $\mathbf{d}({\mathbf{u}}_{k})$ in the input space $\mathcal{I}$. Nevertheless, KPCA has been successfully applied for the design-space assessment of a destroyer-type vessel in [123], where the approximate pre-images technique proposed in [125] is used, and aerodynamic optimization of NACA foil [97], RAE-2822 and CRM wing [126].
Figure 26: Block diagram for the solution of shape optimization problems through AE